A wealthy philanthropist has established the following endowment for a hospital. The details are as follows: a cash deposit of $ 15 M one year from now; an annual cash deposit of $3M per year for the next five years. The first $3M will start today. At the end of 5 years, the hospital will also receive a lump sum payment of $19M. Assuming the cost of money is 4%, what is the value of this endowment in today’s dollars?
Answer #1
1 B C D E F G
2 Time Period Cash InFlow or Income in USD and Million Factors using Return Rate 4% Present Value of Cash InFlow or Income at Return Rate of 4% Cash Outflow or Cost in USD and Million Present Value of Cash Outflow or Cost at Return Rate of 4%
3 0 1 0 3 3
4 1 0.961538462 0 15 14.42307692
5 1 0.961538462 3 2.884615385
6 2 0.924556213 0 3 2.773668639
7 3 0.888996359 0 3 2.666989076
8 4 0.854804191 0 3 2.564412573
9 5 19 0.821927107 15.61661503
10 Total Present Value NA NA 15.61661503 NA 28.3127626
11 Net Present Value or NPV -12.70

The value of this endowment in today’s dollars = USD 12.70 Million (Negative)

Formula

1 B C D E F G
2 Time Period Cash InFlow or Income in USD and Million Factors using Return Rate 4% Present Value of Cash InFlow or Income at Return Rate of 4% Cash Outflow or Cost in USD and Million Present Value of Cash Outflow or Cost at Return Rate of 4%
3 0 =1/(1+0.04)^B3 =C3*D3 3 =F3*D3
4 1 =1/(1+0.04)^B4 =C4*D4 15 =F4*D4
5 1 =1/(1+0.04)^B5 3 =F5*D5
6 2 =1/(1+0.04)^B6 =C6*D6 3 =F6*D6
7 3 =1/(1+0.04)^B7 =C7*D7 3 =F7*D7
8 4 =1/(1+0.04)^B8 =C8*D8 3 =F8*D8
9 5 19 =1/(1+0.04)^B9 =C9*D9
10 Total Present Value NA NA =SUM(E3:E9) NA =SUM(G3:G8)
11 Net Present Value or NPV =E10-G10